OpenStudy (anonymous):

What is the slope and y-intercept of y = -4x + 7?

OpenStudy (anonymous):

i think the slope is -4 and the intercept is 7

OpenStudy (anonymous):

That's the answer I got, I just wanted to check it :)

OpenStudy (anonymous):

you can give me a medal if you are happy.

OpenStudy (anonymous):

Yeah, it's in the form y=mx+c, where m=slope/gradient, and c = y-intercept

OpenStudy (anonymous):

y = mx + c

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

Same thing

OpenStudy (anonymous):

m is the gradient or slope and c is the y intercept

OpenStudy (anonymous):

@iplayffxiv please stop begging for a medal. Many people contributed to this, and it wouldn't be fair if I only gave it to you.

OpenStudy (anonymous):

I wish I could best response all of you. Thank you for the help <3

OpenStudy (anonymous):

y = -4x + 7 y= mx + b ^m is slope ^ b is y intercept.